i need help!!!! way behind on home work and have no idea how to do this algebra! im working on perfect square trinomials. and factoring polynomials. what would 6bsquared+11b+3 how do i solve? and x squared-18x+81=0 how do i solve????
Answer by stanbon(75887)

6bsquared+11b+3 how do i solve?
---
You cannot "solve" it because it is not an equation.
You can factor it:
6b^2 + 11b + 3
Think of two numbers whose product is 6*3 = 18
and whose sum is 11.
The numbers are 9 and 2
Rewrite the problem but replace 11b by 9b+2b
---
6b^2 + 9b + 2b + 3
Factor the 1st two and the last two terms:
3b(2b+3) + (2b+3)
Factor again:
(2b+3)(3b+1)
===========================================
and x squared-18x+81=0
x^2 - 18x + 81
Think of two numbers whose product is 81 and sum is -18.
The numbers are -9 and -9
------------------------------
Rewrite the problem:
x^2 -9x -9x + 81 = 0
x(x-9) -9(x-9) = 0
(x-9)(x-9) = 0
x = 9 with multiplicity 2
==============================
Cheers,
Stan H.
==============================
-----------------